OpenStudy (anonymous):

9th grade word problems 1.Coach Johnson is 10 years older than Tim. When Tim started football 3 years ago, Coach Johnson was twice his age. How old is Tim now? Numerical Answers Expected 2.The sum of three consecutive integers is 267. What is the largest integer? Numerical Answers Expected! 3.A store offers a 15% discount on all items in the store during a sale. All store employees get an additional 10% employee discount. In the equation below, C represents the cost of an item for an employee including tax, and x represents the original price of the item. C = 0.765x + 0.06

OpenStudy (anonymous):

Number 3 continued.... C = 0.765x + 0.06(0.765x) Chris is an employee at the store. He plans to buy a coat during the sale. The price of the coat before the discount is $100. How much will Chris pay for the coat during the sale?

OpenStudy (beginnersmind):

If you show your work it will be easier to give help that's appropriate for your level. Also, consider posting one problem at a time.

OpenStudy (pfenn1):

2. Sum of three consecutive integers could be determined by the equation: n+(n+1)+(n+2)=267 Solve for n

OpenStudy (pfenn1):

1. If t is Tim's age now and T is the Coach's age now, we know T=t+10 T-3=2(t-3) subatitute t+10 into the second equation for T and solve for t

OpenStudy (pfenn1):

3. Are there two different equations for C or is the second equation for C the correct one?